OpenStudy (jskinny0831):

How much money should be invested now (rounded to the nearest cent), called the initial investment, in a Treasury Bond investment that yields 5.75% per year, compounded quarterly for 6 years, if you wish it to be worth $12,000 after 6 years? Hint: A(t) = A0(1 + start fraction r over n end fraction)nt, where A(t) is the final amount, A0 is the initial investment, r is the growth rate expressed as a decimal, n is the number of compounding periods per year, and t is time in years. A. $8,519.54 B. $8,505.65 C. $13,458.78 D. $8,580.23

OpenStudy (jskinny0831):

@soprano.h.d0816 @RhondaSommer @DanJS

OpenStudy (danjs):

yeah this is the same sort of exponential function, y=A*b^x except when you figure the rate of change b on this one, it will only be 1/4 of that total 5.75 yearly percent., that is the r/n thing,

OpenStudy (danjs):

n is the number of compounding cycles each period of t, 1 year

OpenStudy (danjs):

and then instead of just raised to the t power, compounding once per year, it is now n*t, 4*t, since it compounds quarterly, 4 times each year

OpenStudy (jskinny0831):

A. $8,519.54 B. $8,505.65 C. $13,458.78 D. $8,580.23

OpenStudy (jskinny0831):

so its B? @DanJS

OpenStudy (danjs):

this one they give you the desired future value when t=6 years, and ask for the initial investment \[\large A(t) = A _{0}*(1+\frac{ .0575 }{ 4 })^{4*t}\] that is the function, when t=6, you want A(6) = 12000 \[\huge A(6) = 12000= A _{0}*(1+\frac{ .0575 }{ 4 })^{4*t}\]

OpenStudy (danjs):

solve for that initial amount needed Ao, t=6

OpenStudy (jskinny0831):

okay i got a weird number

OpenStudy (danjs):

\[\huge 12000= A _{0}*(1.014375)^{24}\]

OpenStudy (danjs):

12000/(1.014375)^24

OpenStudy (jskinny0831):

oh so its 8519?

OpenStudy (danjs):

yep

OpenStudy (jskinny0831):

okay and one more question if thats okay

OpenStudy (danjs):

sure

OpenStudy (jskinny0831):

The half-life of a substance is how long it takes for half of the substance to decay or become harmless (for certain radioactive materials). The half-life of a substance is 8.6 days and there is an amount equal to 15 grams now. What is the expression for the amount A(t) that remains after t days, and what is the amount of the substance remaining (rounded to the nearest tenth) after 37 days? Hint: The exponential equation for half-life is

OpenStudy (jskinny0831):

A. A(t) = 15(0.5)8.6t, 0.0 gram remaining B. A(t) = 15(0.5)t/8.6, 0.8 gram remaining C. A(t) = 8.6(15)(0.5)t, 0.0 gram remaining D. A(t) = 15(0.5)8.6/t, 12.8 grams remaining

OpenStudy (jskinny0831):

i just need this one so i can go to sleep

OpenStudy (danjs):

this is the same form of general exponential y(t) = A*b^t

OpenStudy (jskinny0831):

so how do i plug it in

OpenStudy (danjs):

The given info is start amount at t=0 is 15g half is left after t=8.6

OpenStudy (jskinny0831):

so 8.6 is C?

OpenStudy (danjs):

each of those is a point on the graph, time t=0 and amojnt y(t) = 15 grams y(0) = 15 * (b)^0 = 15* 1 so A=15, you have y(t) = 15*b^t

OpenStudy (jskinny0831):

A. A(t) = 15(0.5)8.6t, 0.0 gram remaining B. A(t) = 15(0.5)t/8.6, 0.8 gram remaining C. A(t) = 8.6(15)(0.5)t, 0.0 gram remaining D. A(t) = 15(0.5)8.6/t, 12.8 grams remaining

OpenStudy (jskinny0831):

so its gonna be A right

OpenStudy (danjs):

the other thing is half is left after t=8.6, so when half is gone, y(t)/A will be 1/2 or 0.5

OpenStudy (danjs):

1/2 = b^8.6 that gives you the b value,

OpenStudy (jskinny0831):

so the answer is A?

OpenStudy (jskinny0831):

@DanJS

OpenStudy (jskinny0831):

or C?

OpenStudy (jskinny0831):

I'm sorry i just really wanna finish since I'm timed @DanJS

OpenStudy (danjs):

y(t) = 15*(0.922564)^t for t=37 days y(37) = about 0.7 g left

OpenStudy (jskinny0831):

A. A(t) = 15(0.5)8.6t, 0.0 gram remaining B. A(t) = 15(0.5)t/8.6, 0.8 gram remaining C. A(t) = 8.6(15)(0.5)t, 0.0 gram remaining D. A(t) = 15(0.5)8.6/t, 12.8 grams remaining which answer would it be then

OpenStudy (jskinny0831):

i just need to know the letter for the answer @DanJS

OpenStudy (danjs):

prolly B

OpenStudy (danjs):

they just wrote the thing in a different form

OpenStudy (jskinny0831):

you sure?

OpenStudy (danjs):

yes

OpenStudy (jskinny0831):

thank you it was right

OpenStudy (danjs):

yes, just study up that general exponential y=A*b^x, all the probs were really the same, just different context